OpenStudy (anonymous):

prove and justify (tanN + cotN)^2=sec^2N+csc^2 N

OpenStudy (dumbcow):

\[=\tan^{2}N + 2\tan N \cot N +\cot^{2} N\] Note: tan*cot = 1 \[= \tan^{2} N +2 +\cot^{2} N\] from pythagorean identities tan^2 = sec^2 -1 cot^2 = csc^2 -1 \[=(\sec^{2} N -1) +2+(\csc^{2} N -1)\] \[=\sec^{2} N + \csc^{2} N\]